e researcher knows she needs to spin the blood so that it has a centripetalacceleration or 1200g the distance mom the prvot to the blood sample s cm at whatsargenta speed ood spin? he tangential speed that the blood will spin at is m/s
Answer #1

centripetal acceleration

ac=V2/r

tangential speed

V=sqrt(rac)=sqrt(0.3*1200*9.81)

V=59.43 m/s